OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (campbell_st):

add the equations together to eliminate y x - y = 16 2x +y = 2 ---------- 3x = 18 then x = 6 substitute into the 1st equation to find y 6 - y = 16 subtract 6 from both sides -y = 12 multiply by -1 y = -12 the solution is x = 6 y = -12

OpenStudy (anonymous):

ok